Rose is one-third as old as her mother. The sum of their ages is 100.
How old are each of them?
I would really apreciate some help with this I am trying to help my son but it has been long time for me and I am very rusty thank you for your time

Found 2 solutions by stanbon, awang1996:
Answer by stanbon(75887)

Rose is one-third as old as her mother. The sum of their ages is 100.
How old are each of them?
-----------------
Let mother's age be 3x yrs.
Then Rose is x yrs old.
---------
EQUATION:
x + 3x = 100
4x = 100
x = 25 (Rose's age)
3x= 75 (Mother's age)
=================
Cheers,
Stan H.

Answer by awang1996(50)

Rose is one-third as old as her mother. The sum of their ages is 100.
How old are each of them?

3r=m
r+m=100
3r+r=100
4r=100
divide both sides by 4

r=25
m=75